# RV-Syn: Rational and Verifiable Mathematical Reasoning Data Synthesis based on Structured Function Library

### v1

@@ 4.3 Further Analysis
Figure 3: The performance of the model as the training data size increases. The "Average" refers to the mean performance across five datasets: MATH-500, GSM8K, GSM-Hard, College Math, and OlympiadBench.
Scalability Study.
@@
In this section, we study the scalability of our method while maintaining high-quality problem generation. We compare our method with previous approaches by training on scaling datasets using LLaMA-3-8B-Instruct AI@Meta (2024), and further scale our method to 500k to explore its scalability at larger scales. The results in Figure 3 demonstrate a strong scaling curve, with performance consistently improving as the dataset size increases. Our method consistently outperforms other methods, and notably, achieves better performance with only half the dataset size (50k vs. 100k). This advantage is particularly pronounced on challenging or computation-intensive datasets, demonstrating the superiority of our method for enhancing advanced mathematical reasoning capabilities.
Answer Annotation Model Matters.
Dataset Original Re-annotation Performance Change
Jiuzhang3.0 30.5 42.1 +38.0%
MathScale 26.5 41.9 +58.1%
MMIQC 34.9 43.7 +25.2%
NuminaMath 38.7 43.9 +13.4%
Analysis of Problem Quality.
@@
In this part, we quantitatively analyze the difficulty and efficiency of the synthesis problem using three key metrics, following PromptCoT Zhao et al. (2025). First, accuracy serves as an indicator of problem difficulty. Lower performance signifies a higher level of difficulty. For datasets without ground-truth labels, we adopt the approach detailed in PromptCoT Zhao et al. (2025) by leveraging a more powerful model, Qwen2.5-Math-72B-Instruct, to generate reference answers, applying self-consistency with 8 rollouts to enhance reliability. Second, the average number of reasoning tokens provides insights into the complexity of the reasoning process. A larger average number of tokens suggests that the model necessitates more extensive reasoning steps. Finally, the average accuracy on the benchmark after tuning reflects the dataset’s contribution to improving model performance. Larger improvements in benchmark accuracy observed after fine-tuning suggest that the respective dataset is more effective in enhancing the model’s capabilities. Experimental results in Table 4 show a strong consistency among these three metrics, and our method achieves the best results across all three difficulty evaluation metrics.
@@
Table 4: Difficulty and efficacy evaluation for different mathematical datasets. Accuracy: Performance of Qwen2.5-Math-7B-Instruct on the problems in different datasets. Avg. Reasoning Tokens: Average number of tokens in reasoning processes generated by DeepSeek-R1-Distill-Qwen-7B when processing the problems in different datasets. Avg. Benchmark Accuracy: Performance of LLaMA-3-8B-Instruct after fine-tuning on different datasets.
Correctness Verification.
In this part, we investigate the verifiable characteristics of our method to ensure data correctness. We sample 1,000 data points from the training data of each method and evaluate them using GPT-4o. We validate this evaluation method in Appendix B and analyze the error rates for both problems and solutions.
@@
Table 5: Error rate analysis of training data across different methods. The problem error refers to issues such as inconsistency or conflict in problem conditions and statements; the solution error indicates computational mistakes or misinterpretation of the problem in the solution.
As shown in Table 5, firstly, the problem error rate reflects the quality of the synthesized problems. It can be observed that problems generated by humans exhibit a lower error rate, while directly synthesized problems contain more errors. Our method approaches the level of control that humans exert in problem design. Secondly, the solution error rate reflects the extent to which each method controls the correctness of training data. Since other methods cannot naturally obtain the ground truth, their solution error rates are relatively high. Further quality control requires heavy reliance on powerful models for post-filtering, annotation, voting, or scoring, which incurs significant additional costs. In contrast, our method inherently provides the ground truth and enables rule-based matching, ensuring an extremely low error rate. We then manually inspect our data labeled as erroneous and find that most solutions have correct final answers, with only minor errors in intermediate steps.
@@ 5 Conclusion
In this paper, we introduced RV-Syn, a novel approach to mathematical problem synthesis that enhances both rationale and verifiability. It first constructed a graph-format function library, which was composed of a large amount of mathematical functions and was scalable to insert new ones. Then, it generated the new solutions by combining functions into computation graphs in a controllable manner, which can be then executed to support solution verification. Finally, it back-translated these computation graphs to problems. Our experiments demonstrate that RV-Syn outperforms existing methods, achieving higher accuracy with less data. This efficiency highlights its potential for improving the reasoning capabilities of LLMs while minimizing resource consumption. RV-Syn represents a significant advancement in reasoning data generation, offering a scalable and reliable framework. Future work can explore extending it to other reasoning domains and further refining control mechanisms for broader applicability.
@@
Appendix A Function Library Statistic
Here, we show the function library statistic in Table 6 and Table 7.
Topic Proportion
Number Theory 0.204
Prealgebra 0.082
Intermediate Algebra 0.080
Algebra 0.102
Precalculus 0.290
Counting & Probability 0.119
Geometry 0.121
Miscellaneous 0.003
Table 6: Topic proportions in function library. Topics for each function were labeled according to the categories in the MATH dataset.
@@ Appendix A Function Library Statistic
Appendix B Correctness Evaluation Validation
Considering the substantial cost associated with human evaluations, we use one of the most powerful models currently available, GPT-4o, to conduct correctness evaluation. To ensure its validity, we conducted an experiment using GPT-4o to evaluate the accuracy of Qwen2.5-Math-7B-Instruct on the benchmark test set and compared the results with ground-truth evaluations. As shown in Table 8, the evaluation results—74.25% (GPT-4o) vs. 76.61% (Ground truth)—demonstrate a close alignment. False positive samples account for only 1.89% of all cases, while the false negative rate is even higher, indicating that GPT-4o tends to be more stringent in its evaluation. This is likely because it not only verifies the final answer but also checks intermediate steps. Considering that even ground-truth-based evaluations may contain some degree of error, we believe this discrepancy falls within an acceptable range. The prompt we use is shown in Appendix C.
Ground Truth
Correct (+) Incorrect (-) Total
GPT-4o Correct (+) 0.7236 0.0189 0.7425
Incorrect (-) 0.0425 0.2150 0.2575
Total 0.7661 0.2339 1.0000
Table 8: Confusion matrix for GPT-4o evaluation.
Appendix C Prompts
Prompt for Computational Graph Regeneration Fine-tuning Given a set of mathematical operation functions: ``` {FUNCTIONS} ``` Write a ‘main‘ function that intelligently combines these operations to solve a meaningful and non-trivial mathematical problem.
Prompt for Problem Back-translation Given a piece of code that solves a mathematical problem: ``` {CODE} ``` First, understand what the code snippet is doing within the ‘<analysis></analysis>‘ tags. Then, identify the original problem the functions are solving and output it within the ‘<problem></problem>‘ tags. Finally, rewrite the problem concisely, in a manner similar to questions found in textbooks or test papers, and output it within the ‘<final problem></final problem>‘ tags. The question should have only one answer.
Appendix D Examples of function similarity
@@
Appendix E Format Example for Problem Decomposition

### v2

@@ 4.3 Scalability Study
Figure 3: The performance of the model as the training data size increases. The “Average” refers to the mean performance across five datasets: MATH-500, GSM8K, GSM-Hard, College Math, and OlympiadBench.
@@
In this section, we study the scalability of our method. We compare our method with previous approaches by training on scaling datasets up to 300k using LLaMA-3-8B-Instruct AI@Meta (2024). The results in Figure 3 demonstrate a strong scaling curve, with performance consistently improving as the dataset size increases. Our method consistently outperforms other methods, and notably, achieves better performance with only half the dataset size (50k vs. 100k). This advantage is particularly pronounced on challenging or computation-intensive datasets, demonstrating the superiority of our method for enhancing advanced mathematical reasoning capabilities.
4.4 Comparison with Human-Crafted Data
To rigorously benchmark the quality of our synthesized data, we compare RV-Syn against datasets that contain a significant portion of high-quality, human-crafted problems. We include two such strong baselines in our analysis: MMIQC Liu et al. (2024), which contains QA pairs from the Mathematics Stack Exchange, and Numina-Math LI et al. (2024), a large collection of both real-world human-crafted and synthesized math problems, which also serves as the source of our initial seed data. As illustrated in Figure 4, human-crafted datasets such as MMIQC and Numina-Math consistently outperform most synthetic methods, reflecting the superior quality of human-authored data over synthetic data, though they face significant scalability challenges. In contrast, RV-Syn demonstrates a clear competitive advantage: at nearly every data scale, it achieves the highest average performance, even surpassing Numina-Math, the very dataset used as its seed. This result validates that our synthesis process does not merely replicate existing examples, but instead generates novel, high-quality instructions that are more effective for model training than the original seed data itself. Consequently, RV-Syn successfully bridges the gap between scalability and quality, producing synthetic data that rivals human expertise without sacrificing performance. We provide further discussion on data diversity in Appendix D.
Figure 4: Comparison between RV-Syn and human-crafted data. Numina-Math serves as the seed data for our synthetic dataset. Previous state-of-the-art synthetic data methods are indicated by gray lines.
4.5 Analysis of Data Efficiency
Higher complexity through Rational Synthesis.
@@ 4.5 Analysis of Data Efficiency
First, we posit that the solution-first, rational synthesis approach leads to higher-quality and more complex problems. A higher-quality dataset can provide richer supervision signals per example, thus improving training efficiency. To quantify this, we analyze the difficulty and efficacy of different datasets using three key metrics, following PromptCoT Zhao et al. (2025). First, accuracy serves as an indicator of problem difficulty. Lower performance signifies a higher level of difficulty. For datasets without ground-truth labels, we adopt the approach detailed in PromptCoT Zhao et al. (2025) by leveraging a more powerful model, Qwen2.5-Math-72B-Instruct, to generate reference answers, applying self-consistency with 8 rollouts to enhance reliability. Second, the average number of reasoning tokens provides insights into the complexity of the reasoning process. A larger average number of tokens suggests that the model necessitates more extensive reasoning steps. Finally, the average accuracy on the benchmark after tuning reflects the dataset’s contribution to improving model performance. Larger improvements in benchmark accuracy observed after fine-tuning suggest that the respective dataset is more effective in enhancing the model’s capabilities. Experimental results in Table 3 show a strong consistency among these three metrics, and our method achieves the best results across all three difficulty evaluation metrics. Specifically, problems synthesized by RV-Syn are the most difficult (lowest accuracy at 55.6%), require the longest reasoning chains (highest token count), and yield the largest performance improvement after fine-tuning (+34.1%). This demonstrates that our solution-first, rational synthesis approach successfully generates more challenging and effective training instances compared to other methods.
@@
Table 3: Difficulty and efficacy evaluation for different datasets. Accuracy: Performance of Qwen2.5-Math-7B-Instruct on the problems in different datasets. Avg. Reasoning Tokens: Average number of tokens in reasoning processes generated by DeepSeek-R1-Distill-Qwen-7B when processing the problems. Avg. Benchmark Accuracy: Performance of LLaMA-3-8B-Instruct after fine-tuning on different datasets.
Higher Quality through Verifiability.
In this part, we investigate the verifiable characteristics of our method to ensure data correctness. We sample 1,000 data points from the training data of each method, evaluate them and analyze the error rates for both problems and solutions (see Appendix  B for details). As shown in Table 4, firstly, the problem error rate reflects the quality of the synthesized problems. It can be observed that problems generated by humans exhibit a lower error rate, while directly synthesized problems contain more errors. Our method approaches the level of control that humans exert in problem design. Secondly, the solution error rate reflects the extent to which each method controls the correctness of training data. Since other methods cannot naturally obtain the ground truth, their solution error rates are relatively high. Further quality control requires heavy reliance on powerful models for post-filtering, annotation, voting, or scoring, which incurs significant additional costs. In contrast, our method inherently provides the ground truth and enables rule-based matching, ensuring an extremely low error rate. We then manually inspect our data labeled as erroneous and find that most solutions have correct final answers, with only minor errors in intermediate steps.
@@
Table 4: Error rate analysis across different methods. The problem error refers to issues such as inconsistency or conflict in problem conditions and statements; the solution error indicates computational mistakes or misinterpretation of the problem in the solution.
4.6 Ablation Study
To provide a comprehensive understanding of RV-Syn, we conduct ablation studies to analyze the impact of our graph sampling strategies and the controllability of problem difficulty.
Impact of sampling strategies.
We introduce three strategies to sample function nodes for computational graph regeneration: co-occurrence-based, topic-based, and edgeless. To evaluate their individual contributions, we synthesize 20k training samples using each strategy exclusively and train LLaMA-3-8B-Instruct. As shown in Table 5, each strategy excels in different areas. For example, the co-occurrence strategy is strongest on GSM8K. The edgeless strategy performs best on MATH-500. The topic-based strategy shows particular strength on the more complex GSM-Hard and OlympiadBench datasets. This validates our decision to use a mix of all three, creating a balanced and robust dataset that leverages their complementary advantages.
Strategy GSM8K MATH Hard College Olym.
Co-occur. 77.9 35.2 35.3 26.5 11.0
Topic 77.2 35.2 39.2 26.7 12.9
Edgeless 77.3 39.6 37.8 27.1 12.6
Table 5: Ablation study on different sampling strategies.
Controllability of problem difficulty.
A key advantage of RV-Syn is the fine-grained control over synthesized problem complexity. We analyze our ability to control problem complexity by varying the number of function nodes sampled during graph regeneration. We train models on 50k datasets synthesized with a fixed number of nodes (1, 2, and 3 nodes) and evaluate them on MATH-500 and GSM-Hard. As shown in Table 6, models trained on single-node problems achieve the highest performance on MATH-500, a dataset emphasizing direct application of middle-school-level concepts. In contrast, performance on the more calculation-intensive GSM-Hard improves with increasing node count, indicating that multi-node graphs effectively model complex, multi-step reasoning. This demonstrates RV-Syn’s ability to deliberately calibrate difficulty for targeted training needs.
Sampled Nodes MATH-500 GSM-Hard
1 Node 49.0 42.6
2 Nodes 48.6 43.8
3 Nodes 47.2 44.4
Table 6: Ablation on problem difficulty control by varying sampled node counts.
4.7 Analysis of Long-CoT Distillation
@@ 5 Conclusion
In this paper, we introduced RV-Syn, a novel approach to mathematical problem synthesis that enhances both rationale and verifiability. It first constructed a graph-format function library, which was composed of a large number of mathematical functions and was scalable to insert new ones. Then, it generated the new solutions by combining functions into computation graphs in a controllable manner, which can be then executed to support solution verification. Finally, it back-translated these computation graphs to problems. Our experiments demonstrate that RV-Syn outperforms existing methods, achieving higher performance with less data. This efficiency highlights its potential for improving the reasoning capabilities of LLMs while minimizing resource consumption. RV-Syn represents a significant advancement in reasoning data generation, offering a scalable and reliable framework. Future work can explore extending it to other reasoning domains and further refining control mechanisms for broader applicability.
@@
Acknowledgments
This paper was partially supported by the National Natural Science Foundation of China No. 92470205 and Beijing Major Science and Technology Project under Contract no. Z251100008425002. Xin Zhao is the corresponding author.
Appendix A Prompts for Data Generation Pipeline
Prompt for Computational Graph Regeneration Fine-tuning Given a set of mathematical operation functions: ‘‘‘ {FUNCTIONS} ‘‘‘ Write a ‘main‘ function that intelligently combines these operations to solve a meaningful and non-trivial mathematical problem.
Prompt for Problem Back-translation Given a piece of code that solves a mathematical problem: ‘‘‘ {CODE} ‘‘‘ First, understand what the code snippet is doing within the ‘<analysis></analysis>‘ tags. Then, identify the original problem the functions are solving and output it within the ‘<problem></problem>‘ tags. Finally, rewrite the problem concisely, in a manner similar to questions found in textbooks or test papers, and output it within the ‘<final problem></final problem>‘ tags. The question should have only one answer.
Appendix B Details for Correctness Evaluation
Considering the substantial cost associated with human evaluations, we use one of the most powerful models currently available, GPT-4o, to conduct correctness evaluation. To ensure its validity, we conducted an experiment using GPT-4o to evaluate the accuracy of Qwen2.5-Math-7B-Instruct on the benchmark test set and compared the results with ground-truth evaluations. As shown in Table 7, the evaluation results—74.25% (GPT-4o) vs. 76.61% (Ground truth)—demonstrate a close alignment. False positive samples account for only 1.89% of all cases, while the false negative rate is even higher, indicating that GPT-4o tends to be more stringent in its evaluation. This is likely because it not only verifies the final answer but also checks intermediate steps. Considering that even ground-truth-based evaluations may contain some degree of error, we believe this discrepancy falls within an acceptable range. The prompt we use is shown in Appendix A.
Ground Truth
Correct (+) Incorrect (–) Total
GPT-4o Correct (+) 0.7236 0.0189 0.7425
Incorrect (–) 0.0425 0.2150 0.2575
Total 0.7661 0.2339 1.0000
Table 7: Confusion matrix for GPT-4o evaluation.
Appendix C Example of Problem Decomposition
Appendix D Analysis of Data Diversity
A key limitation of data augmentation methods like MetaMath Yu et al. (2024) is their tendency to produce problems that are structurally similar to the original seeds. This often results from the inherent paraphrasing nature of in-context learning, leading to limited diversity. For example, a seed problem asking to “choose 3 captains from a team of 11 people” might be augmented to “choose 2 captains from a team of 9 people”, a trivial variation. RV-Syn also utilizes seed data, but in a fundamentally different manner. Instead of rewriting whole problems, RV-Syn decomposes them to extract a library of atomic mathematical operations (functions). Our synthesis process then generates new problems not by modifying an existing problem, but by creating entirely new computational graphs by combining these atomic functions. This combinatorial approach allows for the creation of novel reasoning paths and complex problem structures that can be fundamentally different from any single seed problem, thereby fostering genuine diversity.
To empirically validate this claim, we quantitatively compare the diversity of data generated by RV-Syn against MetaMath, relative to the seed data. We employed three metrics: $\Delta$ Mean of Embedding Variance: The change in the average variance of problem embeddings. A positive value indicates increased diversity. Centroid Cosine Distance and Maximum Mean Discrepancy (MMD2): Both measure the dissimilarity between the generated data distribution and the original seed distribution. Larger values signify greater divergence.
Method $\Delta$ Mean of Emb. Variance Centroid Cosine Dist. MMD2
MetaMath -0.0017 0.0170 0.000825
RV-Syn +0.0019 0.0398 0.002493
Table 8: Diversity and similarity analysis. RV-Syn increases embedding variance and generates data that is more distinct from the seed set compared to MetaMath.
The results, presented in Table 8, strongly support our hypothesis. MetaMath shows a negative change in embedding variance, suggesting it slightly reduces diversity. In sharp contrast, RV-Syn increases diversity. Furthermore, the significantly larger Centroid Cosine Distance and MMD2 values for RV-Syn confirm that its output diverges more substantially from the seed distribution. This demonstrates that our function-centric, combinatorial synthesis is more effective at producing novel and diverse problems than simple seed-based augmentation.
Appendix E Statistics of the Synthesis Pipeline
Figure 6: Statistical analysis of the RV-Syn pipeline. (a) Distribution of mathematical topics within the function library. (b) Difficulty distribution of the final synthesized problems. (c) Proportions of data filtered out at different stages of the synthesis process.
Here, we present statistics for the function library and the final dataset. Figure 6 (a) shows the topic distribution in our function library, where each function is labeled according to the categories defined in the MATH dataset. Figure 6 (b) displays the difficulty distribution of the synthesized dataset, with the majority of samples concentrated in the higher-difficulty range. This distribution helps explain why RV-Syn achieves more substantial performance gains on more challenging benchmarks. Figure 6 (c) details the filtering proportions during the synthesis pipeline. We observed that 16% of the generated computational graphs failed to execute and were discarded. Among the remaining valid graphs, an additional 37% of the generated problem–solution pairs were filtered out because no Chain-of-Thought solution matched the ground-truth answer produced by graph execution. This automated, graph-based quality control mechanism is crucial for ensuring the high correctness of our final dataset. Table 9 provides additional statistics for the function library.
@@ Appendix E Statistics of the Synthesis Pipeline
Appendix F Impact of the Answer Annotation Model
Dataset Original Re-annotation Performance Change
Jiuzhang3.0 30.5 42.1 +38.0%
MathScale 26.5 41.9 +58.1%
MMIQC 34.9 43.7 +25.2%
Numina-Math 38.7 43.9 +13.4%
Appendix G Influence of Seed Dataset
Since RV-Syn utilizes the high-quality Numina-Math dataset as seed data, we determine whether the performance gains stem from our synthesis framework or merely the superior seed data. We conduct a controlled experiment using the Jiuzhang3.0 baseline. We replace its original web-corpus seed data with Numina-Math and train it on 50k synthesized samples. The results in Table 11 show that simply using better seed data with a direct synthesis method (Jiuzhang3.0) does not improve performance; in fact, the score slightly drops compared to its original setting (39.5 vs. 42.1). We attribute this to the fact that direct synthesis methods rely heavily on the diversity of broad pre-training corpora, and constraining them to a structured QA dataset may limit their generative potential. In contrast, RV-Syn achieves a significantly higher average score (44.92), confirming that our improvements are driven by the rational synthesis framework rather than the seed data alone.
Method Seed Data Avg. Score
Jiuzhang3.0 Original (Web) 42.1
Jiuzhang3.0 Numina-Math 39.5
RV-Syn (Ours) Numina-Math 44.9
Table 11: Impact of seed dataset on performance (50k data, LLaMA-3-8B). Comparing our method against Jiuzhang3.0 seeded with Numina-Math.
@@ Appendix G Influence of Seed Dataset
Appendix H Examples of Function Similarity